# Discourse Heuristics For Paradoxically Moral Self-Correction

### v1

@@ Abstract
Moral self-correction has emerged as a promising approach for aligning the output of Large Language Models (LLMs) with human moral values. However, moral self-correction techniques are subject to two primary paradoxes. First, despite empirical and theoretical evidence to support the effectiveness of self-correction, this LLM capability only operates at a superficial level. Second, while LLMs possess the capability of self-diagnosing immoral aspects of their output, they struggle to identify the cause of this moral inconsistency during their self-correction process. To better understand and address these paradoxes, we analyze the discourse constructions in fine-tuning corpora designed to enhance moral self-correction, uncovering the existence of the heuristics underlying effective constructions. We demonstrate that moral self-correction relies on discourse constructions that reflect heuristic shortcuts, and that the presence of these heuristic shortcuts during self-correction leads to inconsistency when attempting to enhance both self-correction and self-diagnosis capabilities jointly. Based on our findings, we propose a solution to improve moral self-correction by leveraging the heuristics of curated datasets. We also highlight the generalization challenges of this capability, particularly in terms of learning from situated context and model scales. Warning: this paper contains offensive language.
@@
Guangliang Liu [^1] footnotemark: 1   Zimo Qi211footnotemark: 1   Xitong Zhang1   Kristen Marie Johnson1 1Michigan State University  2Johns Hopkins University {liuguan5,zhangxit,kristenj}@msu.edu    zqi15@jh.edu
@@ 1 Introduction
Self-correction is a post-hoc approach that guides LLMs to refine their previous output according to the given instructions Madaan et al. (2023); Kamoi et al. (2024). It has become a popular technique for improving the quality of LLMs’ generations, and its application in enhancing morality, i.e., moral self-correction Ganguli et al. (2023); Liu et al. (2024b, e), effectively reduces harmful and stereotypical content from LLM output.
Prior studies reveal two key paradoxes concerning the effectiveness of moral self-correction within LLMs. Paradox1: While moral self-correction appears effective in enhancing the perceived moral correctness of LLM responses Liu et al. (2024b); Ganguli et al. (2023), this capability remains superficial, as evidenced by limited alterations of hidden states or the requirement of ground-truth answers in instructions Liu et al. (2024c); Huang et al. (2024). Paradox2: There is a lack of consistency between self-diagnosis and self-correction Liu et al. (2024d), suggesting a disconnect between an LLM’s capability of identifying moral issues, e.g., morally unaligned or incorrect output, and addressing them effectively, which can only be done if the LLM knows why the choice it made was morally incorrect.
Furthermore, prior studies have shown that neural language models are capable of generalization across tasks due to internalization of discourse constructions, rather than true language understanding Misra and Mahowald (2024); Chen et al. (2024), and generalization across moral reasoning tasks without understanding true moral pragmatics due to the distributional semantics of LLMs Liu et al. (2025). Additionally, these models exhibit reliance on shallow heuristics (shortcuts) across tasks Dziri et al. (2023); Sun et al. (2024); Yuan et al. (2024).
These findings motivate a plausible hypothesis which we have explored in this work: shallow heuristics in self-correction may enable LLMs to make self-correction decisions without requiring induced immorality in their hidden states or relying on self-diagnosis, thereby addressing both paradoxes. However, this does not mean LLMs make all self-correction decisions based on shallow heuristics. Our analysis builds on recent findings that emphasize construction-based approaches over syntactic rules for studying generalization in LLMs Zhou et al. (2024); Weissweiler et al. (2025); Bunzeck et al. (2025).
@@
In summary, the main contributions of this work are as follows: we address the two paradoxes by identifying novel heuristics underlying self-correction; showcase the potentials and pitfalls of utilizing these heuristics to improve self-correction; and reveal the generalization challenges of self-correction.
@@ 3 Preliminary Study
In this section, we present the experimental setup and findings which demonstrate that LLMs are capable of performing self-correction, without needing jointly accurate self-diagnosis. We also find evidence that specific discourse constructions may be more beneficial than general constructions.
@@
Task Formulation. Table 1 outlines the task formats used for evaluating self-correction and self-diagnosis capabilities. For all stereotype categories, we adopt a consistent self-correction instruction and task format, following prior work Ganguli et al. (2023); Liu et al. (2024c, b). For self-diagnosis, we extend the task format from Schick et al. (2021) by prompting LLMs to assess whether any stereotypes are present toward a social group.
@@ 3.2 General Discourse Construction
In this subsection, we test if a general discourse construction is informative enough to enhance the self-correction and self-diagnosis capabilities. Based on the BBQ Benchmark, the general discourse construction follows the format: Context + biased answer + statement. One example is “We couldn’t start the board meeting at 9 am today because a woman and a man were late. The woman was the secretary. There are gender stereotypes.” This construction explicitly specifies the stereotypes involved, the social groups they target, and the contextual events in which they arise.
@@
Table 3: Experimental Results for Self-correction (top) and Self-diagnosis (bottom) After Fine-tuning With the Proposed General Discourse Construction. The effect of this discourse construction varies with improved cases which are underlined.
Table 3 presents the self-correction and self-diagnosis performance by fine-tuning a Llama3.2-3B model with the corpus reformatted to use our discourse construction. The effects of the general discourse construction varies across capabilities and stereotype categories. Among the six experiments, only half show improvement, two in self-diagnosis and one in self-correction, suggesting the need for more tailored discourse constructions specific to each capability.
@@ 4 Heuristics in Self-correction
To answer these questions, we first propose a discourse construction by referencing the task format associated with each capability. We then conduct an ablation study on its components to identify an effective construction, and finally, we provide empirical evidence that there does exist shallow heuristics underlying those constructions.
@@
Motivation. Previous studies Misra and Mahowald (2024); Ma et al. (2024); Bunzeck et al. (2025) have explored corpus construction at various levels (e.g., sentence, utterance, and discourse) and trained LLMs from scratch to investigate how to allow LLMs to implement language acquisition. Unlike prior research, our target capabilities reflect higher-level social cognition beyond basic language-relevant capabilities***We do not state that LLMs can truly understand language.. To this end, we fine-tune pretrained LLMs using carefully designed discourse constructions, aiming to leverage their existing language proficiency while enabling self-correction and self-diagnosis. We begin by constructing discourse structures that encompass the full range of task-relevant formats, and conduct ablation studies to identify which components are essential.
@@ 4.2 Self-correction
According to Table 4, All-Statement can improve self-correction performance. To further explore this, we refine the Statement to be more situated in order to cover more details and explanations about why the Context conveys social stereotypes. One example is shown within the following textbox. If LLMs still fail in capturing more situated statement, we can confidently claim that LLMs do not rely on Statement for generalization. With the inclusion of the Situated Statement, we expected LLMs to attend to it due to its greater informativeness. However, its introduction unexpectedly degrades self-correction performance, even falling below the baseline, except for the Disability stereotype with the 1B model. This is very strong evidence that LLMs can bypass stereotype awareness to perform self-correction, and intentionally enhancing stereotype awareness may have adverse effects. We believe this is one potential reason that LLMs tend to capture heuristics for moral self-correction.
@@
Situated Context and Event. We are motivated to explore the role of context and events because human language acquisition occurs through situated communication, and pragmatics is acquired within situated context Cicourel (2011); Sap et al. (2022); Beuls and Van Eecke (2024). Table 4 has shown the importance of context; Table 5 shows our additional analysis of the Context and Event. To isolate the effect of the Event, we use the Statement instead of the Action component, as the Action has already been shown to have a significant positive impact when used alone.
@@ 5 Conflicts and Generalization
Given the heuristics proposed for 1B and 3B models, we conduct generalization tests and our experimental results suggest that: (1) conflicts emerge when attempting to enhance both capabilities simultaneously (Section 5.1), and (2) given a certain size of LLMs and our found heuristics, the self-correction performance can be easily improved for both in-domain stereotypes and out-of-domain stereotypes (Section 5.2).
@@ 5.1 Conflicts Between Capabilities
Ideally, we would expect that enhancing one capability could also benefit the other. For example, training LLMs in moral self-correction might implicitly develop their ability to perform self-diagnosis as well. Table 6 presents the self-diagnosis performance when self-correction is enhanced. For the considered models, enhancing self-correction leads to performance improvements only for half of experiments, while it results in performance drops—below baseline—for the rest. Those empirical observations indicate that there does exist conflicts between those two capabilities, and we believe this conflict stems from our finding that the heuristics in self-correction exclude stereotypes awareness, addressing Paradox2.
@@ 5.2 Generalization
1B|3B Age Nation Gender SES Disable Age Nation Gender SES Disable
Baseline .767 .757 .838 .682 .875 .841 .907 .891 .807 .908
Individual .841 .810 .873 .719 .882 .903 .963 .933 .868 .954
Mixed .742 .787 .866 .725 .875 .875 .967 .940 .887 .973
Table 7: In-domain Generalization By Mixing the Fine-tuning Corpus of Five Representative Stereotypes. Individual represents fine-tuning with the discourse within one stereotype in Section 4. Mixed means that we mix the fine-tuning dataset of five stereotypes and test the fine-tuned model on each stereotype. The 3B model shows good in-domain generalization but 1B model does not, implying the generalization of heuristics does rely on model sizes. Please refer to Appendix B for more experimental results for other models.
@@
Table 7 presents the in-domain generalization results across five stereotype categories. Across all models and stereotypes, we observe consistent performance improvements when using the Mixed dataset for the 3B model except the Age stereotype for which Mixed still improve self-correction better than the baseline. However, once the model size decrease to the 1B, Mixed is worse than Individual, except for the SES stereotype. This suggests the capability limitation of small LLMs, which is aligned with previous findings Liu et al. (2024e); Schick et al. (2021); Zhao et al. (2021).
Table 8 presents the experimental results on three out-of-domain stereotypes using the Mixed fine-tuning corpus. Consistent with the in-domain generalization results, the 3B models show improved performance after fine-tuning with the Mixed corpus, whereas the 1B model exhibits limited generalization capability. In summary, we highlight the conflict between improving self-diagnosis and enhancing self-correction, and demonstrate that the identified heuristics exhibit strong generalization in LLMs of certain sizes.
@@
Table 8: Out-of-domain Generalization By Mixing the Fine-tuning Corpus of 5 Representative Stereotypes. 3B model shows good out-of-domain generalization but 1B does not. Please refer to Appendix B for more experimental results for other models.
@@ 8 Limitations
In this paper, we use social stereotype mitigation as a representative task, while noting that other morality-relevant tasks, such as implicit toxicity detection and moral judgment, can also be explored within this framework. Our investigation of self-correction is constrained to the fine-tuning setting due to resource limitations, therefore we can not overlook the impact of pre-training. Training LLMs from scratch to further validate the discourse structures proposed in this study would be more concrete. Intuitively, LLMs should exhibit a degree of language comprehension to perform well on high-level tasks that require social cognition. However, such capabilities remain highly challenging for the NLP community and lie beyond the scope of this paper. Therefore, we refrain from discussing language acquisition and instead focus on morality-relevant capabilities. Our study of interests in this paper is moral self-correction, but it would be great to explore the heuristics underlying self-diagnosis.
@@ B.2 Results of larger models
All-Statement-subaction .986 .993 1.0 .973 1.0 .898 .873 .877 .939 .875
Table 11: Experimental Results on Llama3-8b-it/Mistral-v0.3-7b-it for Self-correction Across Different Discourse Constructions. All includes all possible components: context, statement, action. All - * indicates a component was removed from the setting of All. The optimal performance is highlighted in bold. Across all experiments, only three out of ten All - Statement (Disability in Llama, Age and Nation in Mistral) do harm to the performance compared to the All, suggesting that LLMs do not need stereotype awareness for successful self-correction. Note: We select Llama3-8b due to its architectural consistency with the 1B/3B models in our paper. However, since the BBQ benchmark is widely used, this model has already been fine-tuned on it and achieves near perfect performance. Nevertheless, our framework still demonstrates measurable improvements.
This section presents experiments on larger models, LLaMA-3-8B-Instruct and Mistral-7B-Instruct-v0.3, using LoRA fine-tuning with a rank of 64 and a learning rate of 1e-5. It should be noted that we select Llama3-8b due to its architectural consistency with the 1B/3B models in our study. However, since the BBQ benchmark is widely used, this model has already been fine-tuned on it and achieves near perfect performance. Nevertheless, our framework still demonstrates measurable improvements.
Table 11 summarizes the self-correction performance across all discourse constructions. Compared to smaller models, the 7B/8B parameter models achieve superior performance, with LLaMA-3-8B-Instruct attaining perfect accuracy in four out of five tasks. These results further validate the effectiveness of our proposed constructions for larger-scale LLMs.
Compared with smaller LLMs, there are some important observations. Removing the Context or Action components do harm to the performance relative to the full construction (All). However, only three out of ten All - Statement (specifically, Disability in Llama, Age and Nation in Mistral) underperform the All construction. This suggests that that explicit stereotype awareness is not essential for successful self-correction. These findings align with the shallow heuristics hypothesis presented in Section 4.2 of our paper.
Table 12 present self-diagnosis performance when self-correction is enhanced. There are two key-observations: (1) The overall performance is not higher than smaller models significantly (primarily between 0.6-0.7). (2) The conflicts are slightly mitigated in larger-scale models. The empirical findings necessitate fine-grained analysis of self-diagnosis, which is essential to understand how LLMs perform moral tasks.
Table 13 and Table 14 demonstrate the in-domain and out-of-domain generalization capability of both models. The results show strong generalization of moral self-correction in LLMs for larger-scale models, thereby extending the finding from section 5.2.
@@
Llama3|Mistral SexOrientation Physical Religion SexOrientation Physical Religion
Baseline .947 .959 .920 .759 .849 .785
Mixed .988 .989 .983 .951 .972 983
Table 14: Out-of-domain Generalization of Llama3-8b-it/Mistral-v0.3-7b-it By Mixing the Fine-tuning Corpus of 5 Representative Stereotypes. 7B/8B models show good out-of-domain generalization

### v2

@@ Abstract
Moral self-correction has emerged as a promising approach for aligning the output of Large Language Models (LLMs) with human moral values. However, moral self-correction techniques are subject to two primary paradoxes. First, despite empirical and theoretical evidence to support the effectiveness of self-correction, this LLM capability only operates at a superficial level. Second, while LLMs possess the capability of self-diagnosing immoral aspects of their output, they struggle to identify the cause of this moral inconsistency during their self-correction process. To better understand and address these paradoxes, we analyze the discourse constructions in fine-tuning corpora designed to enhance moral self-correction, uncovering the existence of the heuristics underlying effective constructions. We demonstrate that moral self-correction relies on discourse constructions that reflect heuristic shortcuts, and that the presence of these heuristic shortcuts during self-correction leads to inconsistency when attempting to enhance both self-correction and self-diagnosis capabilities jointly. Building on our findings, we propose a method to strengthen moral self-correction through heuristics extracted from curated datasets, underscoring that its generalization is primarily constrained by situational context. Our code and dataset are publicly available at https://github.com/qzm233/SelfcorrectionHeuristics.
@@
Warning: this paper contains offensive language.
Guangliang Liu1thanks: Equal contribution.   Zimo Qi211footnotemark: 1   Xitong Zhang1   Kristen Marie Johnson1 1Michigan State University  2Johns Hopkins University {liuguan5,zhangxit,kristenj}@msu.edu    zqi15@jh.edu
@@ 1 Introduction
Self-correction is a post-hoc approach that guides LLMs to refine their previous output according to the given instructions Madaan et al. (2023); Kamoi et al. (2024). It has become a popular technique for improving the quality of LLMs’ generations, and its application in enhancing morality, i.e., moral self-correction Ganguli et al. (2023); Liu et al. (2024b, e), effectively mitigate harmful and stereotypical content in LLM outputs.
Prior studies reveal two key paradoxes concerning the effectiveness of moral self-correction within LLMs. Paradox1: While moral self-correction appears effective in enhancing the perceived moral correctness of LLM responses Liu et al. (2024b); Ganguli et al. (2023), this capability remains superficial, as evidenced by limited alterations of hidden states or the requirement of ground-truth answers in instructions Liu et al. (2024c); Huang et al. (2024). Paradox2: There is a lack of consistency between self-diagnosis and self-correction Liu et al. (2024d), suggesting a disconnect between an LLM’s capability of identifying moral issues, e.g., morally unaligned or incorrect output, and addressing them effectively, which can only be done if the LLM knows why the decision it made was morally incorrect.
Furthermore, prior studies have shown that neural language models are capable of generalization across tasks due to internalization of discourse constructions, rather than true language understanding Misra and Mahowald (2024); Chen et al. (2024), and generalization across moral reasoning tasks without understanding true morality due to the distributional semantics of LLMs Liu et al. (2025). Additionally, LLMs exhibit reliance on shallow heuristics (shortcuts) across tasks Dziri et al. (2023); Sun et al. (2024); Yuan et al. (2024).
These findings motivate a plausible hypothesis which we have explored in this work: shallow heuristics***However, this does not mean LLMs make all self-correction decisions based on shallow heuristics. in self-correction may enable LLMs to make self-correction decisions without requiring induced immorality in their hidden states or relying on self-diagnosis, thereby addressing both paradoxes. Our analysis builds on recent findings that emphasize construction-based approaches over syntactic rules for studying generalization in LLMs Zhou et al. (2024); Weissweiler et al. (2025); Bunzeck et al. (2025). Construction-based approaches consider that LLMs achieve generalization not through reliance on syntax, but by leveraging the statistical distributions of phrases.
@@
In summary, the main contributions of this work are as follows: we address the two paradoxes by identifying novel heuristics underlying moral self-correction; showcase the potentials and pitfalls of utilizing these heuristics to improve self-correction; and reveal the generalization challenges of self-correction.
@@ 3 Preliminary Study
In this section, we describe the experimental setup and present findings demonstrating that LLMs are capable of performing self-correction even without fully accurate self-diagnosis. While it is intuitive that LLMs must first identify the moral issues in their outputs before effective self-correction can occur, our results suggest otherwise. Moreover, we provide evidence that certain discourse constructions yield greater benefits for self-correction than more general constructions.
@@
Task Formulation. Table 1 outlines the task formats used for evaluating self-correction and self-diagnosis capabilities. For all stereotype categories, we adopt a consistent self-correction instruction and task format, following prior work Ganguli et al. (2023); Liu et al. (2024c, b). For self-diagnosis, we extend the task format from Schick et al. (2021) by prompting LLMs to assess whether any stereotypes are present toward a social group. Schick et al. (2021) focus on the downstream explicit toxicity implied by social stereotypes in LLMs, whereas our work directly examines the stereotypes themselves. The key distinction between explicit toxicity and social stereotypes is that toxicity can often be identified through rich linguistic cues in the text, whereas social stereotypes operate at the pragmatics level Ma et al. (2025); Liu et al. (2025). Since pragmatics is characterized by context-dependence and implication, we explicitly indicate the social groups for which LLMs should make self-diagnosis decisions.
@@ 3.2 General Discourse Construction
In this subsection, we test if a general discourse construction is informative enough to enhance the self-correction and self-diagnosis capabilities. Based on the BBQ Benchmark, the general discourse construction follows the format: Context + biased answer + statement by referring to Figure 1. One example is “We couldn’t start the board meeting at 9 am today because a woman and a man were late. The woman was the secretary. There are gender stereotypes.” This construction explicitly specifies the stereotypes involved, the social groups they target, and the contextual events in which they arise.
@@
Table 3: Experimental results for self-correction (top) and self-diagnosis (bottom) after fine-tuning with the proposed general discourse construction. The impact of this discourse construction varies across cases, with improved instances indicated by underlining.
Table 3 presents the self-correction and self-diagnosis performance by fine-tuning a Llama3.2-3B model with the corpus reformatted to use our discourse construction. The effects of the general discourse construction varies across capabilities and stereotype categories. Among the six experiments, only half show improvement, two in self-diagnosis and one in self-correction, suggesting the need for more tailored discourse constructions specific to each capabilityPlease note that in this paper, we focus on self-correction rather than self-diagnosis. We leave the investigation of effective discourse constructions for self-diagnosis to future work..
@@ 4 Heuristics in Self-correction
To answer these questions, we first propose a discourse construction by referencing the task format associated with self-correction. We then conduct an ablation study on its components to identify an effective construction, and finally, we provide empirical evidence that there does exist shallow heuristics underlying those constructions.
@@
Motivation. Previous studies Misra and Mahowald (2024); Ma et al. (2024); Bunzeck et al. (2025) have explored corpus construction at various levels (e.g., sentence, utterance, and discourse) and trained LLMs from scratch to investigate how to allow LLMs to implement language acquisition. Unlike prior research, our target capabilities reflect higher-level social cognition beyond basic language-relevant capabilitiesWe do not state that LLMs can truly understand language.. To this end, we fine-tune pretrained LLMs using carefully designed discourse constructions, aiming to leverage their existing language proficiency while enabling self-correction and self-diagnosis. We begin by constructing discourse structures that encompass the full range of task-relevant formats, and conduct ablation studies to identify which components are essential.
@@ 4.2 Heuristics in Self-correction
According to Table 4, All-Statement can improve self-correction performance. To further investigate this, we ablate the role of the abstractive statement by refining the Statement to be more situated, thereby providing richer details and explanations of why the Context conveys social stereotypes. One example is shown within the following textbox. If LLMs still fail in capturing more situated statement, we can confidently claim that LLMs do not rely on Statement for generalization. With the inclusion of the Situated Statement, we expected LLMs to attend to it due to its greater informativeness. However, its introduction unexpectedly degrades self-correction performance, even falling below the baseline, except for the Disability stereotype with the 1B model. This is very strong evidence that LLMs can bypass stereotype awareness to perform self-correction, and intentionally enhancing stereotype awareness may have adverse effects. We believe this explains why LLMs tend to rely on heuristics for moral self-correction, while also suggesting that jointly improving self-diagnosis and self-correction may be infeasible.
@@
Situational Learning. We are motivated to explore the role of context and events because human language acquisition occurs through situated communication, and pragmatics is acquired within situated context Cicourel (2011); Sap et al. (2022); Beuls and Van Eecke (2024). The context and event are situated information for the moral self-correction. Therefore, we validate the isolated effects of context and event, independent of other discourse components. Table 5 shows our additional analysis of the Context and Event. To isolate the effect of the Event, we use the Statement instead of the Action component, as the Action itself has already been shown to have a significant positive impact when used alone.
@@
1B|3B Age Nation Gender SES Disable Age Nation Gender SES Disable
Baseline .767 .757 .838 .682 .875 .841 .907 .891 .807 .908
Individual .841 .810 .873 .719 .882 .903 .963 .933 .868 .954
Mixed .742 .787 .866 .725 .875 .875 .967 .940 .887 .973
8B|7B Age Nation Gender SES Disable Age Nation Gender SES Disable
Baseline .906 .987 .955 .897 .993 .838 .837 .695 .810 .849
Individual .983 1.0 1.0 1.0 .938 .938 .857 .860 .988 .882
Mixed .992 .997 .996 .969 .993 .966 .980 .964 1.0 .967
Table 7: In-domain Generalization By Mixing the Fine-tuning Corpus of Five Representative Stereotypes. We report the in-domain generalization performance across different model scales. Individual represents fine-tuning with the discourse within one stereotype in Section 4. Mixed means that we mix the fine-tuning dataset of five stereotypes and test the fine-tuned model on each stereotype. The 3/7/8B model shows good in-domain generalization but 1B model does not, implying the generalization of heuristics does rely on model sizes.
@@ 5 Conflicts and Generalization
Given the heuristics proposed for moral self-correction, we conduct generalization tests and our experimental results suggest that: (1) conflicts emerge when attempting to enhance both capabilities simultaneously (Section 5.1), and (2) given a certain size of LLMs and our found heuristics, the self-correction performance can be easily improved for both in-domain stereotypes and out-of-domain stereotypes (Section 5.2).
@@ 5.1 Conflicts Between Capabilities
Ideally, we would expect that enhancing one capability could also benefit the other. For example, training LLMs in moral self-correction might implicitly develop their ability to perform self-diagnosis as well. Table 6 presents the self-diagnosis performance when self-correction is enhanced. For the considered models, enhancing self-correction leads to performance improvements only for half of experiments, while it results in performance drops, below baseline, for the rest. Those empirical observations indicate that there does exist conflicts between those two capabilities, and we believe this conflict stems from our finding that the heuristics in self-correction exclude stereotypes awareness, addressing Paradox2.
@@ 5.2 Generalization
Table 7 presents the in-domain generalization results across five stereotype categories. Across all models and stereotypes, we observe consistent performance improvements when using the Mixed dataset for the 3B model except the Age stereotype for which Mixed still improve self-correction better than the baseline. For the 7B model, moral self-correction performance improved when using Mixed, consistently across all stereotype categories. For the 8B model, with the exception of SES stereotypes, Mixed outperforms all others and achieves performance close to that of Individual. However, once the model size decrease to the 1B, Mixed is worse than Individual, except for the SES stereotype. This suggests the capability limitation of small LLMs, which is aligned with previous findings Liu et al. (2024e); Schick et al. (2021); Zhao et al. (2021). Table 8 presents the experimental results on three out-of-domain stereotypes using the Mixed fine-tuning corpus. Consistent with the in-domain generalization results, the 3/7/8B models show improved performance after fine-tuning with the Mixed corpus, whereas the 1B model exhibits limited generalization capability.
In summary, we highlight the conflict between improving self-diagnosis and enhancing self-correction, and demonstrate that the identified heuristics exhibit strong generalization in LLMs of certain sizes.
@@
7B SexOrientation Physical Religion
Baseline .759 .849 .785
Mixed .951 .972 983
8B SexOrientation Physical Religion
Baseline .947 .959 .920
Mixed .988 .989 .983
Table 8: Out-of-domain Generalization By Mixing the Fine-tuning Corpus of 5 Representative Stereotypes. 3/7/8B model shows good out-of-domain generalization but 1B does not. Please refer to Appendix B for more experimental results for other models.
@@ 6 Discussion
Jointly optimizing self-diagnosis and self-correction presents an intriguing challenge, given the inherently statistical nature of LLMs. One promising direction to resolve the conflict between these two capabilities is to enable LLMs acquire pragmatic reasoning for morality Chen and Wang (2025); Liu et al. (2025). Although linguistic research suggests that pragmatic reasoning can be approximated through multi-step semantic inference Bergen et al. (2016), there is still no consensus on how to implement it in practice. Nonetheless, such semantics-driven inference can benefit from the distributional semantics inherent to LLMs.
@@ 8 Limitations
In this paper, we use social stereotype mitigation as a representative task, while noting that other morality-relevant tasks, such as implicit toxicity detection and moral judgment, can also be explored within this framework. Our investigation of self-correction is constrained to the fine-tuning setting due to resource limitations, therefore we can not overlook the impact of pre-training. Training LLMs from scratch to further validate the discourse structures proposed in this study would be more concrete. Intuitively, LLMs should exhibit a degree of language comprehension to perform well on high-level tasks that require social cognition. However, such capabilities remain highly challenging for the NLP community and lie beyond the scope of this paper. Therefore, we refrain from discussing language acquisition and instead focus on morality-relevant capabilities.
@@ B.2 Results of larger models
All-Statement
-subaction
.986 .993 1.0 .973 1.0 .898 .873 .877 .939 .875
Table 11: Experimental Results on Llama3-8b-it/Mistral-v0.3-7b-it for Self-correction Across Different Discourse Constructions. All includes all possible components: context, statement, action. All - * indicates a component was removed from the setting of All. The optimal performance is highlighted in bold. Across all experiments, only three out of ten All - Statement (Disability in Llama, Age and Nation in Mistral) do harm to the performance compared to the All, suggesting that LLMs do not need stereotype awareness for successful self-correction.
@@
This section presents experiments on larger models, LLaMA-3-8B-Instruct and Mistral-7B-Instruct-v0.3, using LoRA fine-tuning with a rank of 64 and a learning rate of 1e-5. It should be noted that we select Llama3-8b due to its architectural consistency with the 1B/3B models in our study. However, since the BBQ benchmark is widely used, this model has already been fine-tuned on it and achieves near perfect performance. Nevertheless, our framework still demonstrates measurable improvements. Table 11 summarizes the self-correction performance across all discourse constructions. Compared to smaller models, the 7B/8B parameter models achieve superior performance, with LLaMA-3-8B-Instruct attaining perfect accuracy in four out of five tasks. These results further validate the effectiveness of our proposed constructions for larger-scale LLMs.
Compared with smaller LLMs, there are some important observations. Removing the Context or Action components do harm to the performance relative to the full construction (All). However, only three out of ten All - Statement (specifically, Disability in Llama, Age and Nation in Mistral) underperform the All construction. This suggests that that explicit stereotype awareness is not essential for successful self-correction. These findings align with the shallow heuristics hypothesis presented in Section 4.2 of our paper. Table 12 present self-diagnosis performance when self-correction is enhanced. There are two key-observations: (1) The overall performance is not higher than smaller models significantly (primarily between 0.6-0.7). (2) The conflicts are slightly mitigated in larger-scale models. The empirical findings necessitate fine-grained analysis of self-diagnosis, which is essential to understand how LLMs perform moral tasks.